OpenStudy (anonymous):

How to solve x^3-7x+6=0 and what is the solution??

OpenStudy (anonymous):

by some miracle this actually factors

OpenStudy (anonymous):

If you factor this expression, you will get all three answers. x^3 - 7x +6 factors to (x-2)(x-1)(x+3). Your answers are x=2, x=1, and x=-3

OpenStudy (anonymous):

1 is a zero by inspection because \[1^3-7\times 1+6=0\] so you know it factors as \[(x-1)\times \text{something}\]

OpenStudy (anonymous):

to find the something you can use synthetic division. or you can just read the corrected answer from jhenahan above.

OpenStudy (anonymous):

Thank you all very much!